### v1

# RAGAR: Retrieval Augment Personalized Image Generation Guided by Recommendation

Abstract
Personalized image generation is crucial for improving the user experience, as it renders reference images into preferred ones according to user visual preferences. Although effective, existing methods face two main issues. First, existing methods treat all items in the user historical sequence equally when extracting user preferences, overlooking the varying semantic similarities between historical items and the reference item. Disproportionately high weights for low-similarity items distort users’ visual preferences for the reference item. Second, existing methods heavily rely on consistency between generated and reference images to optimize the generation, which leads to underfitting user preferences and hinders personalization. To address these issues, we propose Retrieval Augment Personalized Image GenerAtion guided by Recommendation (RAGAR). Our approach uses a retrieval mechanism to assign different weights to historical items according to their similarities to the reference item, thereby extracting more refined users’ visual preferences for the reference item. Then we introduce a novel rank task based on the multi-modal ranking model to optimize the personalization of the generated images instead of forcing depend on consistency. Extensive experiments and human evaluations on three real-world datasets demonstrate that RAGAR achieves significant improvements in both personalization and semantic metrics compared to five baselines.
1 Introduction
Personalized image generation has been widely applied in scenarios such as advertising systems and chat software. It aims to render reference images into preferred ones based on user visual preferences. Existing image generation methods [???] often produce similar outputs for different users with similar input, failing to meet users’ need for personalization. As a personalized generation method, [?] extracts user preferences from user historical sequences by treating each item equally, ignoring the varying semantic similarities between the reference item and historical items, such as semantic differences in categories. This disproportionately amplifies the influence of irrelevant items, causing the extracted user preference to deviate from the true visual preferences on the reference item. Fig. 1(a) shows a comparative example, if a user historical sequence includes electronic products and clothes, and the reference item is ”short sleeves”, an overly simplistic approach might generate a striped T-shirt with the Apple logo—-an outcome that misses the user’s true preferences. By incorporating semantic-based retrieval, more relevant visual preferences, such as ”clothing, sport, pink”, can be captured. Hence, we propose the assumption that filtering the items semantically related to the reference item enhances user preference modeling.
Figure 1: A comparative example shows the effect of retrieval and recommendation by contrasting generated results with and without them.
Another challenge lies in defining and evaluating personalization quantitatively. Current methods rely on either human evaluation or large multi-modal models [??], both of which are resource-intensive. Recent work [?] evaluates and rectifies generated images heavily based on their pixel similarity to the reference image, resulting in limited personalization. Inspired by advances in recommendation systems, which excel in learning and assessing personalized preferences, we propose using multi-modal recommendation models to evaluate generated images, shown in Fig. 1(b). These models provide an efficient and effective way to measure personalization. Moreover, by combining models and metrics from the recommendation domain, we validate in Sec. 2 that: 1) items selected through retrieval are more closely aligned with user preferences than randomly selected items; 2) retrieving items semantically related to the reference item enhances the learning of user preferences compared to irrelevant items.
Hence, the paper proposed the Retrieval Augment Personalized Image GenerAtion guided by Recommendation (RAGAR), comprising three key modules: 1) Retrieval Module contains a correlation unit to calculate the semantic similarity between historical and reference items and a fusion unit to fuse the filtered visual features with dynamic weight. 2) Generation module applies a balance calibrator to combine general user preferences derived from the fine-tuned LLM and the retrieval-augmented user preferences to produce images. 3) Reflection module calculates the loss to balance semantics and personalization, guided by the discriminator.
In summary, the key contributions are as follows:
  • We are the first to emphasize the relationship between historical items and the reference item. The corresponding assumptions are supported by data analysis on real-world datasets in Sec. 2.
  • We propose a novel model RAGAR to generate personalized images. We utilize the retrieval assumption realized by calculating the semantic similarity between items to enhance semantic consistency. Moreover, we design a discriminator to refine and assess the generated images, ensuring higher personalization quality.
  • Experiments on three real-world datasets demonstrate that our proposed method can achieve significant improvements in both personalization and semantic alignment compared to five competing methods.
2 Assumptions Validation
In this section, we validate our assumptions through data analysis on three real-world datasets: POG, ML-latest, and Stickers. [^1] Dataset descriptions are provided in Sec. 5.1. Details can be found in Table 1. To evaluate the effectiveness of the retrieval method, we use the multi-modal recommendation model [?] as the ranking model. We compare three types of user historical sequences: 1) ‘Ret’ contains the top-$k$ items selected using the retrieval method described in Sec. 4.3. 2) ‘Exp-Ret’ (Expanded Retrieval) contains the items ranked between $(k,2k]$ by the retrieval method. 3) ‘Random’ contains $k$ randomly selected items. To assess how well these sequences capture user preferences, we use Recall@10 and NDCG@10 as evaluation metrics.
Figure 2: Assumption validation results across three datasets.
The results shown in Fig. 2. ‘Random’ has the lowest values in metrics, demonstrating that randomly selected items are less effective for preference modeling. Our first assumption holds because both ‘Ret’ and ‘Exp-Ret’ achieve higher metrics, suggesting that items related to the reference items are crucial for modeling user preferences. Moreover, both metrics of ‘Ret’ are higher than those of ‘Exp-Ret’, which implies that the second assumption holds. Therefore, both of our assumptions hold in practical applications, motivating us to retrieve higher ranked items to enhance the user preferences modeling.
3 Related Work
Figure 3: The overview framework of the proposed RAGAR model.
Personalized Image Generation.
Advances in diffusion models (DMs) [?] have enabled high-quality image generation [????]. However, personalized image generation where user preferences guide the generation process, remains underexplored. Methods like [?] and [?] rely on natural language instructions, while [???] combine user interactions and instructions to create personalized advertisements or product images. Recent works bridge DMs and large language models (LLMs) [???], such as [?]. Building on this foundation, [?] incorporates user historical sequences to extract personalized features. However, [?] relies solely on consistency loss for optimization, overfitting the reference image’s features while neglecting personalized preferences. End-to-end multi-modal large language models (MLLMs) [???] show promise in personalizing images from textual instructions often overlook fine-grained preferences. Moreover, widely used evaluation metrics like FID [?], SSIM [?], LPIPS [?], and CLIP score [?] do not align well with human judgment, making them inadequate for assessing personalization. Meanwhile, evaluating personalization manually or using MLLMs [??] is inherently resource-intensive.
Recommendation Systems.
The goal of recommendation systems is to deliver personalized content that meets user preferences. These systems often rely on retrieval [??] and recommendation [???] techniques. Under the multi-modal setting, approaches like [?] integrate textual and visual features for enhanced commonsense reasoning, while [?] extends [?] with image understanding adapters. Specifically, [?] captures user-item relationships across modalities, enabling precise preference modeling for personalized recommendations.
4 Method
In this section, we formulate our task, introduce key notation, and outlines the RAGAR mainstream.
4.1 Task Formulation
We introduce the notation and task formulation for personalized image generation. For the user $u$, let the historical sequence be denoted as $\mathbf{S}_{u}=\{I_{1},I_{2},\cdots,I_{N-1}\}$, and let $I_{N}$ represent the reference item. Personalized image generation aims to produce an image that aligns with the user’s personalized preferences and preserves semantic consistency with $I_{N}$. Under the multi-modal setting, each item $I_{j}$ consists of an image and a text description, denoted as $\{img_{j},txt_{j}\}$.
4.2 RAGAR
Our proposed method, RAGAR, is illustrated in Fig. 3. To reduce the impact of irrelevant items, we introduce the retrieval module that calculates semantic similarity scores between the reference item and historical items. Then, the module fuses the visual features of the items with the weight corresponding to the scores to produce retrieval-augmented preference features. Next, the generation module employs a fine-tuned LLM to derive general preference features from multi-modal historical sequences, corrected by the retrieval-augmented preference. A diffusion-based generator then generate images conditioned by corrected preference. Finally, we design a reflection module to evaluate the generated images, balancing personalization and semantics to ensure high-quality results.
4.3 Retrieval Module
The retrieval module aims to integrate visual features of historical items that are semantically relevant to the reference item, including the correlation unit and the fusion unit.
Correlation Unit.
The correlation unit transforms image information from historical items into textual features to compute semantic similarity scores with the reference item. Given the historical sequence $\mathbf{S}_{u}$ and reference item $I_{N}$ for the user $u$, we extract the textual description from the item images using a caption model (e.g., BLIP-2 [?]):
$cap_{i}=\textbf{Caption}(img_{i}),i=0,1,\cdots,N$ (1)
where $cap_{i}$ represents the caption of item $I_{i}$. Next, the textual description is processed into high-dimensional semantic features with a text encoder, such as CLIP [?]:
$\mathbf{E}^{sem}_{i}=\mathbf{Encoder}(cap_{i})$ (2)
where $\mathbf{E}^{sem}_{i}$ denotes the high-level semantic representation of each caption. Then, we calculate the semantic similarity between historical items and the reference item using cosine similarity:
$s_{i}=cosine(\mathbf{E}^{sem}_{i},\mathbf{E}^{sem}_{N}),i=0,1,\cdots,N-1$ (3)
where $s_{i}$ represents the similarity score for each historical item relative to the reference item. The similarity scores are sorted in descending order, and the top-$k$ items form the retrieval sequence $\mathbf{S}^{ret}_{u}$, capturing the most relevant items in the user’s history.
Fusion Unit.
The fusion unit integrates preference features from the retrieval sequence by computing a weighted sum of visual features based on their semantic similarity scores, emphasizing high-similarity items while filtering out low-similarity ones. For the retrieval sequence $\mathbf{S}^{ret}_{u}$, we extract the visual features using a visual encoder:
$\mathbf{E}^{vis}_{i}=\mathbf{Encoder_{V}}(img_{i}),i=0,1,\cdots,N-1$ (4)
where $\mathbf{E}^{vis}_{i}$ denotes the high-level visual representation of each image. Then, the retrieval-augmented preference feature is computed as a weighted sum of these visual features and their corresponding similarity scores:
$\mathbf{P}^{ret}=\sum_{i=1}^{k}\mathbf{E}^{vis}_{i}\cdot s^{\prime}_{i}$ (5)
where $\mathbf{P}^{ret}$ represents the retrieval-augmented preference feature for the user $u$, and $s_{i}^{\prime}$ denotes normalized similarity scores.
4.4 Generation Module
The generation module bridges general user preferences with retrieval-augmented preferences to produce personalized images. It operates in three stages: 1) extracting general preference features using an LLM, 2) correcting the general preference features with the balance calibrator, and 3) generating personalized images using a diffusion-based generator.
Prompt Construction.
To capture user preferences from interactions, we transform interacted items into structured textual descriptions suitable for LLM analysis. Specifically, we adopt the approach from [?] that summarizes items with concise keywords to enhance interpretability and reduce extraneous information. We construct a prompt $p_{k}$ (detailed in supplementary material) populated with item captions and text to extract concise keywords with an LLM, denoted as $\mathbf{LLM}_{k}$:
$w_{i}=\mathbf{LLM}_{k}(p_{k}(cap_{i},txt_{i})),i=0,1,\cdots,N-1$ (6)
where $w_{i}$ is the keyword list for item $I_{i}$. We filter out low-frequency keywords and retain the top-$n$ keywords as $w^{\prime}$. Next, we construct a prompt $p_{g}$ to capture the general user preference. To enhance the expressivity of the LLM, denoted as $\mathbf{LLM}_{g}$, for image generation, we expand its vocabulary by incorporating $L$ additional special tokens ${\left[IMG\{i\}\right]}_{i=1}^{L}$. The tokens and keywords $w^{\prime}$ are then populated to $p_{g}$ (detailed in supplementary material) to extract preferences.
General preference Extraction.
Given the prompt $p_{g}$, the output of $\mathbf{LLM}_{g}$ can be divided into two embeddings:
$[\mathbf{E}^{txt};\mathbf{E}^{img}]=\mathbf{LLM}_{g}(p_{g})$ (7)
where $\mathbf{E}^{txt}$ represents text-related embedding and $\mathbf{E}^{img}\in\mathbb{R}^{L\times d}$ encodes image-related features. To bridge the gap between textual description and visual representation in LLMs, we adopt the Modal Mapper to align $E^{img}$ with the image space following [?], a 4-layer encoder-decoder transformer with trainable queries and two linear layers. Then we concatenate the text-related embedding and the image-related embedding and obtain the multi-modal feature.
$\mathbf{E}^{m}=[\mathbf{E}^{txt};\mathbf{ModalMapper}(\mathbf{E}^{img})]$ (8)
To maximize keyword utilization, we use the text encoder to generate keyword feature $\mathbf{E}^{key}$. We then employ a cross-attention layer to integrate $\mathbf{E}^{m}$ with $\mathbf{E}^{key}$:
$\mathbf{P}^{gen}=[Attn(\mathbf{E}^{m},\mathbf{E}^{key},\mathbf{E}^{key});\mathbf{E}^{key}]$ (9)
where $Attn$ denotes the cross attention layer to fuse features. In this way, we obtain the general preference feature $P^{gen}$.
Balance Calibrator.
As mentioned in Sec. 4.3, the retrieval-augmented preference focuses primarily on item features associated with the reference item. To narrow the gap between the retrieval-augmented preference feature $\mathbf{P}^{ret}$ and the global preference feature $\mathbf{P}^{gen}$, we minimize the calibrator loss between them:
$\mathcal{L}_{\mathrm{cal}}=\|\mathbf{P}^{gen}-\mathbf{P}^{ret}\|_{2}^{2}$ (10)
This adjustment ensures that $P^{gen}$ reflects both the general and retrieval-augmented preferences.
Image Generation.
With the corrected preference, we utilize the diffusion-based generator to generate images. During training, we sample $r$ random noises $\{\epsilon_{t}\}_{t=1}^{r}\sim\mathcal{N}(\mu,\sigma^{2}I)$ and apply reparameterization trick to enable gradient backpropagation through the sampling process:
$\mathbf{z_{t}}=\mu+\sigma\odot\epsilon_{t}$ (11)
The generator then produces $r$ images using the policy gradient update strategy discussed in Sec. 4.5. Additionally, the keyword feature $\mathbf{E}^{key}$ is used to generate a reference-like image, which serves as input to the ranking model for comparison with the personalized images during the reward computation.
$\begin{split}v^{key}&=Unet(\mathbf{E}^{key})\\ v^{gen}_{t}&=Unet(\mathbf{P}^{gen}+\mathbf{z_{t}}),t=0,1,...,r\end{split}$ (12)
where $v^{key}$ represents the image generated using the keyword features, and $v^{gen}$ represents the image generated using the general preference features.
4.5 Reflection Module
As the diffusion-based generator, propagating gradients backwards from the generated images is challenging [?]. To address this issue, we design a two-part reflection module that balances personalization and semantics in the generated images, ensuring both user alignment and semantic consistency.
Personalization Reflection
To provide personalized feedback on generated images, we leverage a pre-trained multi-modal ranking model (RM) to evaluate images, reducing the reliance on expensive manual labeling. Given a set of images $\{v^{ref},v^{key},v^{gen}_{t}\}$, where $v^{ref}$ represents the reference image. By substituting visual features from this set into the sequence of original items, RM assigns the scores and rankings for each item:
$\{\rho,rk\}=\textbf{RM}(S_{u},v)$ (13)
where $\rho\in\{\rho^{ref},\rho^{key},\rho^{gen}_{t}\}$ is the rank score, and $rk\in\{1,2,3\}$ denotes the rank derived from the score. Intuitively, images that align better with the user’s personalization preference are expected to achieve higher scores and lower ranks. We then employ a reward function to guide the model toward generating images that reflect personalization preferences. Inspired by the policy gradient method [?], we accumulate rewards for each noise $\epsilon_{t}$:
$\mathcal{R}_{t}=\sum_{c\in\{ref,key\}}\max\left(0,\rho^{c}-\rho^{gen}_{t}\right)+\delta$ (14)
$\mathcal{L}_{rank}=-\frac{1}{r}\sum_{t=1}^{r}\log_{\pi_{\theta}}(\mathbf{P}^{gen}+\epsilon_{t})\cdot\mathcal{R}_{t}$ (15)
where $\delta$ is a margin that encourages the model to generate higher-quality images, $\log_{\pi_{\theta}}$ computes probability density of the generated embedding under the Gaussian distribution.
Semantic Reflection
To enhance the semantic consistency, we minimize the distance between the mapper’s output in the generation module, $\mathbf{E}^{m}$, and the semantic features of the reference image, $\mathbf{E}^{sem}_{N}$ using the semantic loss:
$\mathcal{L}_{\mathrm{sem}}=\|\mathbf{E}^{m}-\mathbf{E}^{sem}_{N}\|_{2}^{2}$ (16)
Combined with the calibrator loss, generated images align both semantically and visually with the reference.
Joint Reflection
We employ LoRA [?] to update a limited set of parameters in LLM, the Modal Mapper and the attention fusion components. The overall loss is defined as:
$\mathcal{L}=\alpha\cdot\mathcal{L}_{cal}+\beta\cdot\mathcal{L}_{\mathrm{rank}}+\gamma\cdot\mathcal{L}_{\mathrm{sem}}$ (17)
where $\alpha$, $\beta$ and $\gamma$ are adjustable weighting hyper-parameters.
Features POG ML-latest SER30K
#Users 1000 4689 2230
#Items 19242 9742 30739
Table 1: The characteristics of experimental datasets.
5 Experimentation
We evaluate RAGAR on three different scenarios: commodity, movie poster and sticker. Our experiments aim to answer the following research questions:
  • RQ 1: How does RAGAR perform compared to other generative methods in quantitative evaluation metrics?
  • RQ 2: Dose RAGAR outperform baseline methods in human evaluation?
  • RQ 3: What is the impact of each module of RAGAR on performance? Specifically, how do the retrieval module and the reflection module contribute in the generation?
  • RQ 4: Can personalized images generated by RAGAR improve the performance in recommendation systems?
5.1 Experimental Settings
Datasets.
We utilize three real-world datasets to generate personalized images across different scenarios: POG [^2] https://github.com/wenyuer/POG.,ML-latest [^3] https://grouplens.org/datasets/movielens.,SER30K [^4] https://github.com/nku-shengzheliu/SER30K.. POG is a multi-modal dataset of fashion clothing with user interaction history. Due to the large size, we randomly select a subset of users and interactions. ML-latest is a benchmark movie dataset with user ratings. We collect corresponding posters from IMDB [^5] https://www.imdb.com. and split user historical sequences to ensure each contains 20 interactions. SER30K is a large-scale sticker dataset where each sticker is categorized by theme and annotated with an associated emotion label. We merge two random themes into user historical sequences and randomly select a reference sticker from other themes. Dataset statistics are summarized in Table 1.
Datasets Methods Personalization Semantic Alignment
[…]
POG GLIDE -54.02 13.90 59.01 49.01 17.24 17.52 59.17 58.35 27.14
SD-v1.5 -25.88 15.02 62.97 55.46 12.53 24.18 63.72 63.50 12.39
TI -25.64 15.64 63.79 53.68 15.67 24.14 67.83 62.40 15.23
LaVIT -51.09 15.40 63.46 48.77 22.65 24.07 72.04 56.42 24.79
PMG -22.96 14.82 56.74 55.76 5.19 18.48 57.00 63.89 5.02
RAGAR -19.77 15.79 66.88 55.93 17.99 24.28 74.55 59.29 19.08
ML-latest GLIDE -3.20 13.27 29.52 60.34 17.99 18.21 42.31 61.88 26.22
SD-v1.5 -1.97 12.85 31.07 59.74 14.02 18.38 53.59 58.58 14.51
TI -2.58 14.42 32.85 59.33 14.29 25.25 54.17 59.35 14.79
LaVIT -1.54 14.80 34.78 57.01 18.16 21.04 57.56 56.35 19.64
PMG -0.11 15.07 41.80 53.83 7.34 14.30 43.85 53.51 7.29
RAGAR 0.01 16.20 43.45 56.56 15.14 19.04 53.24 56.30 15.75
Sticker GLIDE -0.86 12.27 49.52 59.78 18.00 16.47 48.97 67.88 28.23
SD-v1.5 -0.93 12.48 51.08 59.45 17.38 18.46 49.61 68.41 16.43
TI -0.83 13.24 50.85 59.48 16.76 17.26 50.14 67.17 15.72
LaVIT -0.75 12.64 53.05 57.82 24.41 21.12 58.99 62.27 24.43
PMG -0.73 13.31 50.97 59.93 4.94 18.72 50.92 68.81 4.73
RAGAR 0.99 14.97 55.25 56.71 25.61 17.69 64.41 61.38 27.48
Table 2: Quantitative performance comparison on three datasets in terms of personalization and semantic alignment. The best performance is bold while the second-best is underlined.
Comparison Methods.
We compare RAGAR with five generative baselines, including three DM-based models Glide [?], SD [?] and TI [?], and two LLM-based models LaVIT [?] and PMG [?].
Evaluation Metrics.
We compare RAGAR with baselines through the following quantitative metrics: To evaluate personalization, we calculate the CLIP Personalization Score (CPS) and the CLIP Personalization Image Score (CPIS), which measure the similarity between the generated images and the text description or images representing user preferences, respectively. In addition, we calculate the LPIPS [?] and SSIM [?] to quantify the perceptual similarity. Furthermore, we compute the rank change $\mathbf{\Delta R}$ between the original and generated images. This metric emphasizes improvements for higher-ranked items by diminishing the weight for lower-ranked ones. A larger $\mathbf{\Delta R}$ indicates greater personalization in the image.
$\mathbf{\Delta R}=\frac{rk_{ori}-rk_{gen}}{1+rk_{ori}}$
where $rk$ denoted the rank determined by the ranking model. To evaluate semantic alignment, we calculate CLIP Score (CS), CLIP Image Score (CIS), LPIPS [?] and SSIM [?] for reference images.
Parameter Settings.
To make a fair comparison, all baselines are tuned with a fixed learning rate of $1e^{-5}$ and stable diffusion 1.5 is used as the image generator. For RAGAR, we set the learning rate at $1e^{-5}$. The number of retrieval items is fixed to 5. The number of noise is set at 3. All experiments are conducted on a single NVIDIA-A100 GPU.
5.2 Performance Comparison (RQ1)
The comparison between RAGAR and baselines is shown in Tab. 2. We can draw the following conclusions. First, traditional DM-based methods perform poorly on personalization metrics. Specifically, Glide and SD are the worst performers across five personalization metrics in all datasets. This is because both methods rely on the CLIP model to generate images, which struggles to capture deeper associations between texts. TI introduces a word embedding approach based on SD to capture user preferences. The stylized word embedding significantly improves personalization performance across all datasets (e.g., for $\Delta R$: -25.64 vs. -54.02, -2.58 vs. -3.20, -0.83 vs. -0.86). In terms of semantic alignment, all three methods perform well, and Glide achieves optimal results on SSIM across all datasets. This indicates that traditional methods tend to focus on reconstructing reference images.
Second, LLM-based methods demonstrate better performance, benefiting from their ability to understand and analyze text comprehensively. LaVIT, a multi-modal vision-language transformer, is designed to align textual and visual features effectively. Although it outperforms DM-based methods, it still falls short in personalization ($\Delta R$: -51.09 in the POG dataset). PMG, an LLM-based personalized image generation method, leverages LLMs to extract user preferences. This approach significantly enhances personalization performance, achieving optimal results compared to DM-based baselines. However, it heavily relies on the consistency with reference images, limited the personalization performance.
Third, our proposed method, RAGAR, achieves state-of-the-art (SOTA) performance due to the retrieval for sequence and the rank model for training across the three datasets. The performance boost is more pronounced on the ML dataset compared to POG and Sticker, likely because the amount of data in the front is significantly larger than that in the back. Compared to PMG, the second-best model, our method enhances consistency by semantically retrieving sequences and introducing evaluation metrics that better align with human judgment, resulting in stronger personalization performance.
Methods POG ML-latest Sticker
Per. Sen. Per. Sen. Per. Sen.
ORI 2.26 / 2.16 / 2.18 /
PMG 2.08 1.62 2.1 1.94 2.24 1.88
RAGAR 1.66 1.46 1.74 1.66 1.58 1.12
Table 3: Human evaluation comparison on both personalization and semantic alignment.
5.3 Human Evaluation (RQ2)
To assess the personalization and semantics of RAGAR in real scenarios, we conduct human evaluation to compare it with PMG and original images. We invited 50 volunteers and set up two types of sorting tasks, with 50 cases for each dataset, covering personalization and semantics. For personalization, given 5 historical images along with the generated images, participants are asked to sort images in descending order of likelihood that the user would click them, based on historical click data. For semantics, given the reference image and generated images, participants are asked to sort them in descending order of how well they convey the same meaning as the reference image. As shown in Tab. 3, RAGAR outperforms the baselines on both metrics. This indicates RAGAR better reflects user preference while preserving the semantics.
Figure 4: Ablation study of the retrieval module and the rank rewards.
Figure 5: Examples of image generation across three datasets.
5.4 Ablation Study (RQ3)
We study the effectiveness of key components of RAGAR in the POG dataset, including retrieval and reflection module.
Effect of retrieval module.
We assess the impact of retrieval module on improving semantic consistency and capturing user preferences by excluding it during training. Fig. 4(a) shows that: 1) noisy items in the historical sequence diminish the model’s ability to capture user preferences and interfere with semantics; 2) the retrieval module selects items relevant to user preferences, thereby improving the performance. Furthermore, we investigate the effect of varying retrieval number $k$ during training. The results reveal that $k=5$ yields the best results when historical sequence length is 20. At $k=0$ (no retrieval), performance is poor, while increasing $k$ initially improves performance but declines beyond $k=5$ due to reintroduced noise and redundancy.
Effect of rank rewards.
We investigate the role of rank rewards by excluding it while retaining the other loss functions during training, denoted as ‘w/o’. The results in Fig. 4(b) demonstrate that the model’s performance decreases without rank rewards. Adding the amount $r$ of sample noise $\epsilon$ boosts the model’s performance. Specifically, $r=5$ strikes the optimal balance between personalization and semantic alignment.
5.5 Case Study
We present cases of personalized images generated by RAGAR and PMG across three datasets: POG, ML-latest, and Sticker. Fig. 5 shows eight historical images, a reference image, and generated images for context. The orange dashed boxes highlight the relevant items retrieved by the RAGAR retrieval module, arranged together for clearer visualization. In clothing scenario, RAGAR generates red-and-blue patterned shoes, effectively capturing the user’s preference for shoes while filtering out irrelevant items, such as patent leather shoes, thereby outperforming PMG in focusing on user preferences. In the movie poster scenario, RAGAR reflects preferences for multiple characters appearing together and avoids irrelevant anime aesthetics, which PMG fails to exclude. In the sticker scenario, RAGAR maintains semantic coherence between a cat and a guitar while aligning with the user’s preference for pink ears and a brown cat style. It successfully filters out realistic style preferences, demonstrating its ability to balance personalization and semantics. Additional cases are in the supplementary material.
Methods POG ML-latest
R@10 N@10 R@10 N@10
ORI 0.1765 0.1544 0.3229 0.2653
PMG 0.1804 0.1550 0.3213 0.2682
RAGAR 0.1927 0.1668 0.3299 0.2882
Table 4: Quantitative comparison of the recommendation performances.
5.6 Auxiliary Generation(RQ4)
Generating items beyond existing data can uncover user interests outside current data distribution [?]. Personalized images generated by RAGAR not only excel in display quality but also enhance recommendation tasks. We validate this using the multi-modal recommendation model MICRO[?]. We conduct experiments on two datasets with real interactions: POG and ML-latest. During training of rank model, reference images are replaced with those generated by RAGAR, PMG and original images ORI, respectively. The results in Tab. 4 reveal that generating images beyond original data distribution improves accuracy of the recommendation model. Both PMG and RAGAR outperform ORI across all metrics, demonstrating the benefit of generative methods. Notably, RAGAR shows a stronger capability in modeling user preferences compared to PMG, achieving significant improvements in metrics across both datasets, 6.8% (Recall@10) and 7.6% (NDCG@10) in POG, 2.67% (Recall@10) and 7.4% (NDCG@10) in ML-latest.
6 Conclusions and Future Work
In this paper, we proposed RAGAR for personalized image generation that balances personalization and semantics. With retrieval-augmented preferences, multi-modal alignment, and reflection-based optimization, RAGAR generates high-quality images that enhance both user experience and downstream recommendation performance. Extensive experiments demonstrate our method outperforms five baselines. For future work, we intend to unify preferences and noise in generation process to further enhance personalization.